Google Corporation has no debt on its balance sheet in 2008, but paid $1.6 billion in taxes. Assume that Google’s marginal tax rate is 35% and Google’s borrowing cost is 7%. Assume that investors hold Google stock in retirement accounts that are free from personal taxes. If Google were to issue sufficient debt to reduce its taxes by $1 billion per year permanently, then the amount that Google needs to borrow is?
Answer #1
Interest to be paid to save $1 billion = 1/0.35 = $       2.86 billion
Borrowings to be made = 2.86/7% = $    40.86 billion